OpenStudy (anonymous):

\[f(x)= \frac{3}{1-x^{4}}\]

OpenStudy (anonymous):

ok so I need to find a power series rep'in the function

OpenStudy (anonymous):

\[\sum c_{n}(x-a)^{n}\] ?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}\]

OpenStudy (anonymous):

Just a suggestion, what do you think about doing the Taylor serie for that function? Then you see the different terms and you can make the power serie, uhm

OpenStudy (anonymous):

Taylor and MacLaurin series is the next session...I haven't learned that yet

OpenStudy (anonymous):

ok teach me, how would you use the Taylor series?

OpenStudy (anonymous):

Using geometric series and the formula for the sum, tell me if this helpful

OpenStudy (anonymous):

"Using geometric series and the formula for the sum..." I"m not sure

OpenStudy (anonymous):

neither do I lol I'm going to compute the serie using this method

OpenStudy (anonymous):

lol...sounds good

OpenStudy (anonymous):

Well I'm not sure, but \[a/1-r=\sum_{0}^{n} ar^n\] in the function f(x)=3/1-x^4 a=3; r=x^4 so the series should be \[\sum_{0}^{n} 3(x^4)^n\] Anyway I don't know if this is correct or not, just a suggestion, ask someone else, sorry MathSofiya

OpenStudy (turingtest):

does it say you need to to the Taylor series about x=0, or some other number?

OpenStudy (anonymous):

no this pre-Taylor series...let me get my book

OpenStudy (anonymous):

this section is titled power series

OpenStudy (anonymous):

i'll write more details from this sections in a second, I'll be right back...sorry (10 mins)

OpenStudy (turingtest):

You know, I was just brushing up on this stuff myself; I'm a little rusty on this way of finding power series Perhaps if I can't parse it out for you right now these notes will help http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeriesandFunctions.aspx

OpenStudy (anonymous):

ok i'm back

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} c_{n} x^{n}\] ok so a power series is of this form...lets see ...

OpenStudy (anonymous):

I am beyond silly...wrong chapter...I'M SOO SORRY

OpenStudy (anonymous):

ok the section is titled Representations of functions as power series for f(x)=1/(1+x^2) we have \[\sum_{n=0}^{\infty}(-x^{2})^{n}\]

OpenStudy (anonymous):

\[\frac{1}{1+x^{2}}= \frac{1}{1-(-x^{2})}\]= \[\sum_{n=0}^{\infty}(-x^{2})^{n}\]

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

\[ \frac{1}{1-x}=\sum_{k=0}^{\infty}x^k\] for\(-1<x<1\) replace \(x\) by \(x^4\) radius of convergence does not change

OpenStudy (anonymous):

similarly \[\frac{1}{1+x^2}=\frac{1}{1-(-x^2)}\] replace \(x\) by \(-x^2\) for your power series

OpenStudy (anonymous):

gimmick is often trying to match up with a geometric series

OpenStudy (anonymous):

\[f(x)= \frac{3}{1-x^{4}}\] so here I would have \[f(x)= \frac{3}{1-(x^{4})}\]

OpenStudy (anonymous):

parentheses not really necessary here

OpenStudy (anonymous):

\[\frac{1}{1-x^{4}}=\sum_{k=0}^{\infty}x^4\]

OpenStudy (anonymous):

it is necessary here \[\frac{1}{1+x}=\frac{1}{1-(-x)}\] to make it match up to the geometric series,

OpenStudy (anonymous):

yes what you write is correct

OpenStudy (anonymous):

what should we do with the three in the numerator?

OpenStudy (anonymous):

slap it next to the x?

OpenStudy (anonymous):

it is a constant, pull it out front

OpenStudy (anonymous):

\[\frac{3}{1-x^{4}}=3\sum_{k=0}^{\infty}x^4\]

OpenStudy (anonymous):

\[\frac{1}{1-x^{4}}=\sum_{k=0}^{\infty}x^4\] \[\frac{3}{1-x^{4}}=3\sum_{k=0}^{\infty}x^4=\sum_{k=0}^{\infty}3x^4\]

OpenStudy (zarkon):

where is your k?

OpenStudy (anonymous):

oops that what i get for pasting

OpenStudy (anonymous):

lolz...never copy and paste what I wrote...

OpenStudy (anonymous):

\[\frac{1}{1-x}=\sum_{k=0}^{\infty}x^k\] \[\frac{1}{1-x^4}=\sum_{k=0}^{\infty}x^{4k}\]

OpenStudy (anonymous):

I don't quite see how the left side of the equal sign equals the right side...I copied the first example from the book so it must be right...but could you explain why they're equal

OpenStudy (anonymous):

Ok, so the first was right